# Gazing at Rewards: Eye Movements as a Lens into Human and AI Decision-Making in Hybrid Visual Foraging

### v2

@@ Abstract
Imagine searching a collection of coins for quarters ($0.25$), dimes ($0.10$), nickels ($0.05$), and pennies ($0.01$)—a hybrid foraging task where observers look for multiple instances of multiple target types. In such tasks, how do target values and their prevalence influence foraging and eye movement behaviors (e.g., should you prioritize rare quarters or common nickels)? To explore this, we conducted human psychophysics experiments, revealing that humans are proficient reward foragers. Their eye fixations are drawn to regions with higher average rewards, fixation durations are longer on more valuable targets, and their cumulative rewards exceed chance, approaching the upper bound of optimal foragers. To probe these decision-making processes of humans, we developed a transformer-based Visual Forager (VF) model trained via reinforcement learning. Our VF model takes a series of targets, their corresponding values, and the search image as inputs, processes the images using foveated vision, and produces a sequence of eye movements along with decisions on whether to collect each fixated item. Our model outperforms all baselines, achieves cumulative rewards comparable to those of humans, and approximates human foraging behavior in eye movements and foraging biases within time-limited environments. Furthermore, stress tests on out-of-distribution tasks with novel targets, unseen values, and varying set sizes demonstrate the VF model’s effective generalization. Our work offers valuable insights into the relationship between eye movements and decision-making, with our model serving as a powerful tool for further exploration of this connection. All data, code, and models will be made publicly available.
@@ 2 Related Work
Computational models for goal-directed visual search. Eye movement in visual search is typically guided by five primary factors [133]: bottom-up saliency [63, 121, 132], top-down feature guidance [33, 81, 29], scene properties [12, 54, 120], prior history [129, 31, 67], and item values [4, 79, 3]. A range of computational models have been proposed to model human eye movements in each of these aspects during visual search [42, 125, 57, 74, 126, 21, 22, 13, 71, 124, 115, 16, 91, 16, 65, 64, 45, 2, 83, 142, 49]. Recent deep-learning models, such as [90, 138, 48, 9, 142, 49, 143, 30], have demonstrated success in searching for a single target in complex, naturalistic scenes. However, these models fail to look for multiple instances of multiple target types based on varying values and prevalence. To close this gap, we introduce Visual Forager (VF), capable of searching for multiple instances across different target types. Moreover, while models like GazeFormer and IRL rely on human data for training and struggle to generalize to unseen targets or unseen combinations of values and prevalence in out-of-distribution scenarios, VF can flexibly adapt its foraging strategy under these conditions.
@@ 3.1 Human psychophysics experiments
We conducted an in-lab psychophysics experiment for hybrid visual foraging, schematically illustrated in Fig. 2. The items were randomly arranged on a 16×16 grid, containing either 90, 105, or 120 items, with 20% to 30% of the squares filled with target instances. In each foraging trial, subjects searched for $N\in\{1,2,4\}$ target objects, each with a varying number of target instances. Targets and distractors in the hybrid foraging search arrays were randomly selected from a pool of 2,400 unique items in [14]. A total of 15 subjects were recruited, yielding 750 trials, containing 50514 eye fixations and 12851 mouse clicks. All the experiments are conducted with the subjects’ informed consent and according to the protocols approved by the Institutional Review Board of our institution. See Appendix LABEL:{sec:humanpsyExp} for more details.
@@ 4.1 Target feature modulation
Receptive field sizes in the visual cortex increase progressively across brain areas [43]. This scaling is modeled in current visual recognition systems through eccentricity-dependent pooling operations [49]. Unlike uniform sampling in standard max pooling, receptive field size in eccentricity-dependent pooling layers grows with increasing distance from the current fixation $F_{t}$. Following [49], we replace VGG16’s standard max pooling layers $l_{10}$, $l_{14}$, and $l_{18}$ with eccentricity-dependent pooling [49]. This modified feature extractor produces feature maps $\phi_{S,t}$ for the search image $\mathcal{I}_{S}$ from the last convolutional layer, given the fixation location $F_{t}$. For simplification of mathematical notations, we omit the subscript $t$ in $\phi_{S,t}$ for the rest of the text, unless specified. We use pre-trained ImageNet weights for VGG16 and freeze them for semantic feature extraction. Since pooling layers lack learnable weights, freezing the pre-trained weights of eccentricity-dependent VGG16 does not affect feature extraction quality. Note that we do not use eccentricity-dependent VGG16 to process feature maps of input target images $\mathcal{I}_{T}^{1:N}$ as these target images are often small and they can be processed in high resolution within the foveated region.
@@ 4.4 Training Details
In the first training stage, we focus solely on training VF’s fixation policy $\pi_{f}(\cdot|s)$, omitting the click policy $\pi_{c}(\cdot|s)$. To encourage VF to fixate on high-value targets, we introduce intrinsic rewards aligned with the rewards for correct clicks: VF receives a reward equal to the target’s value for accurate fixations and a penalty of -1 for fixating on distractors. Additionally, a small penalty of -0.01 discourages fixation on blank areas. To further ease the task, we temporarily disable VF’s eccentricity-dependent vision, allowing it full-resolution access to the search image.
@@ 4.5 Baselines and Evaluation Metrics
We propose two evaluation metrics. First, Normalized Score (Norm.Score) refers to the cumulative rewards as a function of the number of clicks within a foraging trial, normalized by the maximum score of UpperBound for that trial. The normalized score has an upper limit of 1, and a higher normalized score indicates that AI models or humans are more effective at making decisions to optimize cumulative rewards. Second, Click Bias Ratio (CBR) reflects the clicking biases of humans and AI models during foraging. We define the proportion of selections of target objects as Proportion Picked (PP) and the proportion of target objects left on the array as Proportion On-Screen (POS). CBR is then calculated as the normalized relative difference between PP and POS: $CBR=\frac{PP-POS}{PP+POS}$. CBR can change over the course of clicks within a trial. If neither humans nor AI models exhibit clicking preferences, PP will equal POS, resulting in a CBR of 0. If either agent prefers one target object over others, PP will be greater than POS, yielding a positive CBR. If an AI model aligns with human clicking bias patterns in a given experiment, their CBR values will share the same signs.
@@ 5.1 Humans and AI models are reward-seeking
Humans and AI models are proficient foragers. We present the Norm.Score of humans and AI models under three conditions where target prevalence, value, or both vary, as shown in Fig. 4. Human subjects achieved Norm. Scores of 87.4%, 84.1%, and 93.1% across these conditions, significantly exceeding chance levels. Notably, even at the first mouse click, the Norm. Score was already well above chance. Similarly, all competitive baselines and our VF model consistently outperformed chance across all click numbers. These results suggest that both human subjects and AI models are effective at hybrid visual foraging, with their mouse click decisions strongly influenced by the values of target objects. However, humans never reached the upper bound of performance, indicating they are not perfect global optimizers. This may be due to imperfect object recognition [83], limited memory capacity [131], and foveated vision with restricted receptive fields [43].
Our VF model outperform all the baseline models. The Norm. Scores of our VF model are 72.6%, 67.1%, and 81.6% across the three conditions, surpassing all baseline models. This indicates that our VF model effectively learns to optimize decision-making in foraging tasks and adapts well to variations in value and prevalence. Among the baseline models, FeatOnly achieved the highest Norm. Score, although it still performed worse than VF. This highlights the importance of visual feature modulation in foraging tasks, but also reveals that feature modulation alone is insufficient for optimal decision-making. While both AvgVal and MaxVal models incorporate value-based guidance into their decision strategies, their lower performance compared to FeatOnly suggests that simply relying on explicit values is ineffective. Rather, value and feature modulations must interact in a more sophisticated manner to achieve optimal performance.
@@ 5.3 Behavioral alignment between humans and VF
Both humans and AI models tend to overpick high-valued targets and underpick low-valued targets. In this analysis, we examine the items selected by humans and AI models from the search arrays and report their click bias ratios (CBR) averaged over all click numbers under the UnValEqPre and UnValUnPre conditions, as shown in Fig. 5A. Consistent with [135], we find that humans tend to overpick the highest-valued targets, as indicated by a positive CBR, while they underpick the lowest-valued targets, reflected by a negative CBR. Interestingly, our VF exhibits similar signs and magnitudes of CBR on these targets, suggesting that it displays analogous click biases based on target values as humans. However, both humans and VF do not exclusively select the highest-valued targets, as their Proportion Picked (PP) for less-valued targets is not exactly zero. This suggests that both humans and VF occasionally prioritize lower-valued targets during foraging (see Fig. S5 in Appendix). Moreover, we also noted a slight discrepancy in the signs of CBR when humans and VF select the second-highest-valued items. This indicates that VF is less sensitive to medium rewards compared to humans (also see Appendix Fig. S5).
VF approximates the saccade size distributions of humans, without training on human eye movements. Human saccade sizes are restricted by physiological limitations of the eye muscles and the necessity for accurate visual processing in the foveal region [43]. We aggregate all saccade sizes from humans and our VF model under 3 conditions and plot their distributions in Fig. 5B. Despite lacking prior training on human eye movements, VF yields a mean saccade size of 4.06 degree , closely approximating the mean saccade size of 4.05 degree for humans. Additionally, to investigate the mechanisms in VF that constrain saccade sizes, we present the distribution of saccade sizes with eccentricity-dependent layers replaced with standard max-pooling layers in the visual feature extractor. As expected, and in line with [49], the saccade sizes increase, suggesting that smaller saccade sizes are partially due to foveal processing.
@@ 5.4 VF generalizes to OOD conditions
5.5 Ablations reveal critical component designs
Ablations UnVal UnVal EqVal
EqPre UnPre UnPre
Behavior Clone 61.7 48.5 60.1
VF (2D-CNN) 75.3 63.7 70.0
Explicit Val. Emb. 69.2 56.7 61.8
W/o Augmentation 51.3 52.0 52.2
Full VF (ours) 72.6 67.1 81.6
Table 1: Ablation studies reveal critical design choices of our VF model. Norm.Score for various ablated models are reported over UnValEqPre, UnValUnPre, and EqValUnPre conditions. See Sec. 5.5 for ablated models. Best is in bold.
We systematically ablated several essential components in our VF model and reported their results in Tab. 1. (1) Rather than using reinforcement learning, we train VF on human eye movements and mouse clicks through supervised learning (Behavior Cloning). The lower Norm. Score of Behavior Cloning indicates that human eye movement data is limited, leading to model overfitting. Hence, the model struggles to generalize to unseen target value and prevalence combinations. (2) We replace the transformer-based decision-making module with a 2D-CNN, referred to as VF(2D-CNN). The lower Norm. Score of this ablated model indicates that the transformer architecture, with its ability to capture long-range dependencies and global context through self-attention, leads to better decision-making. (3) We ablate VF by replacing the learnable value encoder with explicit value embeddings and directly feeding them into the transformer (Explicit Val. Emb.). The small drop in Norm. Score suggests that a learnable value embedding is more effective for making better decisions. (4) We remove the permutations of target and value pairs (W/o Augmentation), resulting in a significant drop in Norm. Score, especially under the EqValUnPre condition. This indicates that the data augmentation in VF is crucial for enhancing generalization to OOD hybrid foraging tasks.
@@ 6 Discussion
In hybrid visual foraging, humans are proficient foragers, directing attention to regions with significantly higher rewards than chance. When there is an imbalance in target prevalence or values, humans tend to over-exploit the most prevalent or high-valued target types. To explain the characteristics of human behavior in hybrid visual foraging, we propose a transformer-based Visual Forager (VF) trained in reinforcement learning. Its cumulative rewards match human performance and surpass other baseline models. Despite zero training on any human data, VF closely approximates human foraging behaviors, including foraging biases and eye movements. Remarkably, VF demonstrates exceptional generalization abilities, flexibly adjusting its foraging strategies to experimental conditions it has never encountered. Our work paves the way for several new research directions in psychology, neuroscience, and AI. We discuss these future works in Appendix Sec. S4.
@@
In this formulation, the first term inside the $\min$ operator is the standard policy gradient objective, while the second term applies the clipping mechanism to ensure that the policy update does not result in excessively large changes. This clipping mechanism is crucial for maintaining the stability of the learning process.
@@
S3.1 Human motor response
@@
S3.2 Human fixation duration
@@
S3.3 Average reward within fixation area
@@
First, we observed that humans occasionally clicked on items they were not directly fixating on, while VF assumes eye movements always align with the locations at which foraging decisions are made. Second, a strong priming effect was evident in humans, especially when target values were equal, showing the long-lasting influence of prior experiences on human decisions. Our VF currently lacks the ability to model such long-term dependencies, as it does not have a working memory integrating reinforcements from past actions into current decisions. Third, in hybrid foraging, humans actively compare fixated items with those in memory, a process known as memory search. Our VF assumes perfect memory search, where all targets are compared to the fixated item simultaneously. Lastly, real-world environments may present additional challenges, such as target occlusions and physical constraints imposed by scene contexts. Extending the study of hybrid visual foraging beyond simplistic stimuli in controlled experimental settings remains an intriguing research direction.

### v3

@@ Abstract
Imagine searching a collection of coins for quarters ($0.25$), dimes ($0.10$), nickels ($0.05$), and pennies ($0.01$)—a hybrid foraging task where observers look for multiple instances of multiple target types. In such tasks, how do target values and their prevalence influence foraging and eye movement behaviors (e.g., should you prioritize rare quarters or common nickels)? To explore this, we conducted human psychophysics experiments, revealing that humans are proficient reward foragers. Their eye fixations are drawn to regions with higher average rewards, fixation durations are longer on more valuable targets, and their cumulative rewards exceed chance, approaching the upper bound of optimal foragers. To probe these decision-making processes of humans, we developed a transformer-based Visual Forager (VF) model trained via reinforcement learning. Our VF model takes a series of targets, their corresponding values, and the search image as inputs, processes the images using foveated vision, and produces a sequence of eye movements along with decisions on whether to collect each fixated item. Our model outperforms all baselines, achieves cumulative rewards comparable to those of humans, and approximates human foraging behavior in eye movements and foraging biases within time-limited environments. Furthermore, stress tests on out-of-distribution tasks with novel targets, unseen values, and varying set sizes demonstrate the VF model’s effective generalization. Our work offers valuable insights into the relationship between eye movements and decision-making, with our model serving as a powerful tool for further exploration of this connection. All data, code, and models are available at https://github.com/ZhangLab-DeepNeuroCogLab/visual-forager.
@@ 2 Related Work
Computational models for goal-directed visual search. Eye movement in visual search is typically guided by five primary factors [136]: bottom-up saliency [66, 124, 135], top-down feature guidance [34, 84, 30], scene properties [12, 57, 123], prior history [132, 32, 70], and item values [4, 82, 3]. A range of computational models have been proposed to model human eye movements in each of these aspects during visual search [44, 128, 60, 77, 129, 21, 22, 13, 74, 127, 118, 16, 94, 68, 67, 47, 2, 86, 145, 51]. Recent deep learning models, such as [93, 141, 50, 40, 145, 51, 146, 31], have demonstrated success in searching for a single target in complex, naturalistic scenes by predicting a sequence of eye fixations and their durations. However, these models fail to look for multiple instances of multiple target types based on varying values and prevalence. To close this gap, we introduce Visual Forager (VF), capable of searching for multiple instances across different target types. Moreover, while models like GazeFormer [93], OAT [40], and IRL [141], rely on human data for training and struggle to generalize to unseen targets or unseen combinations of values and prevalence in out-of-distribution scenarios, VF can flexibly adapt its foraging strategy under these conditions.
@@ 3.1 Human psychophysics experiments
We conducted an in-lab psychophysics experiment for hybrid visual foraging, schematically illustrated in Fig. 2. The items were randomly arranged on a 16×16 grid, containing either 90, 105, or 120 items, with 20% to 30% of the squares filled with target instances. In each foraging trial, subjects searched for $N\in\{1,2,4\}$ target objects, each with a varying number of target instances. Targets and distractors in the search arrays were randomly selected from a pool of 2,400 unique items in [14]. A total of 15 subjects were recruited, yielding 750 trials, containing 50514 eye fixations and 12851 mouse clicks. All the experiments are conducted with the subjects’ informed consent and according to the protocols approved by the Institutional Review Board of our institution. See Appendix LABEL:{sec:humanpsyExp} for more details.
@@ 4.1 Target feature modulation
Receptive field sizes in the visual cortex increase progressively across brain areas [45]. This scaling is modeled in current visual recognition systems through eccentricity-dependent pooling operations [51]. Unlike uniform sampling in standard max pooling, receptive field size in eccentricity-dependent pooling layers grows with increasing distance from the current fixation $F_{t}$. Following [51], we replace VGG16’s standard max pooling layers $l_{10}$, $l_{14}$, and $l_{18}$ with eccentricity-dependent pooling. We replicated the eccentricity-dependent pooling from [51], aligning it with neurophysiological recordings in macaque monkeys. See Sec. S1.4 for the detailed implementation of eccentricity-dependent pooling. This modified feature extractor produces feature maps $\phi_{S,t}$ for the search image $\mathcal{I}_{S}$ from the last convolutional layer, given the fixation location $F_{t}$. For simplification of mathematical notations, we omit the subscript $t$ in $\phi_{S,t}$ for the rest of the text, unless specified. We use pre-trained ImageNet weights for VGG16 and freeze them for semantic feature extraction. Since pooling layers lack learnable weights, freezing the pre-trained weights of eccentricity-dependent VGG16 does not affect feature extraction quality. Note that we do not use eccentricity-dependent VGG16 to process feature maps of input target images $\mathcal{I}_{T}^{1:N}$ as these target images are often small and they can be processed in high resolution within the foveated region.
@@ 4.4 Training Details
In the first training stage, we focus solely on training VF’s fixation policy $\pi_{f}(\cdot|s)$, omitting the click policy $\pi_{c}(\cdot|s)$. VF receives a reward equal to the target value for accurate fixations and a penalty of -1 for fixations on distractors. This is consistent with the rewards and penalty setups for humans. Moreover, clicking on blank areas wastes time and reduces available clicks; thus, we also apply an intrinsic penalty of -0.01 to discourage suboptimal behaviors in VF. See Tab. S5 for the impact of this intrinsic reward via ablation studies. To further ease the task, we temporarily disable VF’s eccentricity-dependent vision, allowing it full-resolution access to the search image.
@@ 4.5 Baselines and Evaluation Metrics
We also include four more baseline models from the existing literature: IVSN [145] and its variant with value modulation and behavioral cloning, along with the pre-trained Gazeformer [93] and Gazeformer fine-tuned on our in-domain data. Since these models do not utilize foveated vision, direct comparisons with our VF model are not entirely equitable. Nevertheless, we include their results to provide a performance reference in Tab. S6.
We propose two evaluation metrics. First, Normalized Score (Norm.Score) refers to the cumulative rewards as a function of the number of clicks within a foraging trial, normalized by the maximum score of UpperBound for that trial. The normalized score has an upper limit of 1, and a higher normalized score indicates that AI models or humans are more effective at making decisions to optimize cumulative rewards. Second, Click Bias Ratio (CBR) reflects the clicking biases of humans and AI models during foraging. We define the proportion of selections of target objects as Proportion Picked (PP) and the proportion of target objects left on the array as Proportion On-Screen (POS). CBR is then calculated as the normalized relative difference between PP and POS: $CBR=\frac{PP-POS}{PP+POS}$. CBR can change over the course of clicks within a trial. If neither humans nor AI models exhibit clicking preferences, PP will equal POS, resulting in a CBR of 0. If either agent prefers one target object over others, PP will be greater than POS, yielding a positive CBR. If an AI model aligns with human clicking bias patterns in a given experiment, their CBR values will share the same signs. We also used scanpath similarity metrics, including ScanMatch [25] and Fixation Edit Distance (FED) [93], to compare the spatial and temporal dynamics of fixation sequences between humans and AI models. See Tab. S4 for the results and discussion on scanpath similarity.
@@ 5.1 Humans and AI models are reward-seeking
Humans and AI models are proficient foragers. We present the Norm.Score of humans and AI models under three conditions where target prevalence, value, or both vary, as shown in Fig. 4. Human subjects achieved Norm. Scores of 87.4%, 84.1%, and 93.1% across these conditions, significantly exceeding chance levels. Notably, even at the first mouse click, the Norm. Score was already well above chance. Similarly, all competitive baselines and our VF model consistently outperformed chance across all click numbers. These results suggest that both human subjects and AI models are effective at hybrid visual foraging, with their mouse click decisions strongly influenced by the values of target objects. However, humans never reached the upper bound performance, indicating they are not perfect global optimizers. This may be due to imperfect object recognition [86], limited memory capacity [134], and foveated vision with restricted receptive fields [45]. We also present human and model scanpath visualizations in Fig. S6. These examples further support our analysis.
Our VF model outperform all the baseline models. The Norm. Scores of our VF model are 72.6%, 67.1%, and 81.6% across the three conditions, surpassing all baseline models. This indicates that our VF model effectively learns to optimize decision-making in foraging tasks and adapts well to variations in value and prevalence. Among the baseline models, FeatOnly achieved the highest Norm. Score, although it still underperformed VF. This highlights the importance of visual feature modulation in foraging tasks, but also reveals that feature modulation alone is insufficient for optimal decision-making. While both AvgVal and MaxVal models incorporate value-based guidance into their decision strategies, their lower performance compared to FeatOnly suggests that simply relying on explicit values is ineffective. Rather, value and feature modulations must interact in a more sophisticated manner to achieve optimal performance.
@@ 5.3 Behavioral alignment between humans and VF
Both humans and AI models tend to overpick high-valued targets and underpick low-valued targets. In this analysis, we examine the items selected by humans and AI models from the search arrays and report their click bias ratios (CBR) averaged over all click numbers under the UnValEqPre and UnValUnPre conditions, as shown in Fig. 5A. Consistent with [138], we find that humans tend to overpick the highest-valued targets, as indicated by a positive CBR, while they underpick the lowest-valued targets, reflected by a negative CBR. Interestingly, our VF exhibits similar signs and magnitudes of CBR on these targets, suggesting that it displays analogous click biases based on target values as humans. However, both humans and VF do not exclusively select the highest-valued targets, as their Proportion Picked (PP) for less-valued targets is not exactly zero. This suggests that both humans and VF occasionally select lower-valued targets during foraging. Moreover, we also noted a slight discrepancy in the signs of CBR when humans and VF select the second-highest-valued items. This indicates that VF is less sensitive to medium rewards compared to humans (Fig. S5).
VF approximates the saccade size distributions of humans, without training on human eye movements. Human saccade sizes are restricted by physiological limitations of the eye muscles and oculomotor controls [45]. We aggregate all saccade sizes from humans and our VF model under 3 conditions and plot their distributions in Fig. 5B. Despite lacking prior training on human eye movements, VF yields a mean saccade size of 4.06 degree , closely approximating the mean saccade size of 4.05 degree for humans. Additionally, to investigate the mechanisms that constrain saccade sizes in VF, we present the distribution of saccade sizes with eccentricity-dependent layers replaced with standard max-pooling layers in the visual feature extractor. In line with [51], the saccade sizes increase, suggesting that smaller saccade sizes are partially due to foveal processing.
@@ 6 Discussion
In hybrid visual foraging, humans are proficient foragers, directing attention to regions with significantly higher rewards than chance. When there is an imbalance in target prevalence or values, humans tend to over-exploit the most prevalent or high-valued target types. To explain the characteristics of human behavior in hybrid visual foraging, we propose a transformer-based Visual Forager (VF) trained in reinforcement learning. We designed experiments using 2D grid worlds which provide precise control over experimental variables, minimize confounding factors, and make the analysis of cognitive processes more focused. VF’s cumulative rewards match human performance and surpass other baseline models. Despite zero training on any human data, VF closely approximates human foraging behaviors, including foraging biases and eye movements. Remarkably, VF demonstrates exceptional generalization abilities, flexibly adjusting its foraging strategies to experimental conditions it has never encountered. We also clarify that novel stimuli for humans do not equate to an OOD task for AI since humans accumulate years of visual experiences which enables them to generalize effectively. However, we emphasize that while humans exhibit remarkable generalization based on prior experience and flexible cognition, AI models are still inferior to humans from our results. Our work paves the way for several new research directions in psychology, neuroscience, and AI. We discuss these future works in Appendix Sec. S4.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its NRFF award NRF-NRFF15-2023-0001 and its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025). We also acknowledge Mengmi Zhang’s Start Up Grant from Agency for Science, Technology, and Research (A*STAR), and Start Up Grant from Nanyang Technological University.
@@ Acknowledgements
S1.4 Eccentricity dependent pooling
We replicated the eccentricity dependent pooling from [51], aligning it with neurophysiological recordings in macaque monkeys. A conversion of 30 pixels to 1 degree of visual angle (dva) was applied to match human behavioral experiments. To examine how the layer-specific scaling factor $\gamma_{l}$ affects search efficiency and average saccade amplitudes, we varied $\gamma_{l}$ by a coefficient $\beta$. Tab. S1 shows that increasing $\beta$ reduces saccade size and cumulative rewards in UnValEqPre. VF with $\beta=1$ best matches human saccades and achieves comparable rewards.
Humans $\beta=1$ $\beta=2$ $\beta=4$
Avg.Sac.Size (dva) 4.05 4.06 2.26 2.12
NormScore (%) 87.4 72.6 6.4 2.95
Table S1: Ablation of layer-specific scaling factor in eccentricity dependent pooling
@@
In this formulation, the first term inside the $\min$ operator is the standard policy gradient objective, while the second term applies the clipping mechanism to ensure that the policy update does not result in excessively large changes. This clipping mechanism is crucial for maintaining the stability of the learning process. Our VF leverages Generalized Advantage Estimation (GAE) [108] for advantage calculation and TD($\lambda$) for value estimation [11]. This choice is motivated by the computational efficiency of TD($\lambda$) compared to Monte Carlo sampling [56], as noted in the work of GAE.
@@
S3.1 Ablations reveal critical component designs
Ablations UnVal UnVal EqVal
EqPre UnPre UnPre
Behavior Clone 61.7 48.5 60.1
VF (2D-CNN) 75.3 63.7 70.0
Explicit Val. Emb. 69.2 56.7 61.8
W/o Augmentation 51.3 52.0 52.2
Full VF (ours) 72.6 67.1 81.6
Table S3: Ablation studies reveal critical design choices of our VF model. Norm.Score for various ablated models are reported over UnValEqPre, UnValUnPre, and EqValUnPre conditions. See Sec. S3.1 for ablated models. Best is in bold.
We systematically ablated several essential components in our VF model and reported their results in Tab. S3. (1) Rather than using reinforcement learning, we train VF on human eye movements and mouse clicks through supervised learning (Behavior Cloning). The lower Norm. Score of Behavior Cloning indicates that human eye movement data is limited, leading to model overfitting. Hence, the model struggles to generalize to unseen target value and prevalence combinations. (2) We replace the transformer-based decision-making module with a 2D-CNN, referred to as VF(2D-CNN). The lower Norm. Score of this ablated model indicates that the transformer architecture, with its ability to capture long-range dependencies and global context through self-attention, leads to better decision-making. (3) We ablate VF by replacing the learnable value encoder with explicit value embeddings and directly feeding them into the transformer (Explicit Val. Emb.). The small drop in Norm. Score suggests that a learnable value embedding is more effective for making better decisions. (4) We remove the permutations of target and value pairs (W/o Augmentation), resulting in a significant drop in Norm. Score, especially under the EqValUnPre condition. This indicates that the data augmentation in VF is crucial for enhancing generalization to OOD hybrid foraging tasks.
S3.2 Human motor response
@@
S3.3 Human fixation duration
@@
S3.4 Average reward within fixation area
@@
S3.6 Scanpath similarity
We calculated the results of ScanMatch and Fixation Edit Distance (FED) to assess the scanpath similarity between humans and AI models (Tab. S4). Best are in bold. Results show that our VF generates more human-like scanpaths than chance but with lower similarity than within-subject scanpaths across repeated trials and between subjects performing the same trials.
Within-subject Between-subject VF Chance
ScanMatch $\uparrow$ 0.3556 0.3173 0.2474 0.1596
FED $\downarrow$ 39.60 47.50 55.90 60.10
Table S4: Scanpath similarity.
S3.7 Penalty ablation
The $-1$ penalty for distractors ensures consistent rewards for both the agent and humans. Clicking on blank areas wastes time and reduces available clicks, so we apply a $-0.01$ intrinsic penalty to discourage this and prevent suboptimal behavior in VF. We conduct ablation experiments to assess the impact of this instinctive reward and find that it has no significant effect on the final result (see Tab. S5).
-1 -2 -0.01 0
NormScore (%) 75.49 76.5 72.6 79.9
Table S5: NormScore of UnValEqPre condition trained with different instinctive reward.
S3.8 External baseline
We introduced four external baselines that do not utilize foveated vision: (1) IVSN: This baseline iteratively selects the maximum from four-channel similarity maps and applies infinite inhibition of return. (2) IVSN-NN: A variant of IVSN, where the attention map is modulated by value, incorporating an additional neural network module trained via behavior cloning. (3) pre-GF: The pre-trained GazeFormer model [93]. (4) GF: The GazeFormer model fine-tuned on our in-domain data. We tested these models in our OOD tasks. Our VF outperforms all of them (see Tab. S6). The inferior performance of GF and pre-GF compared to our VF is due to their failure to account for descriptive texts of multiple targets with different values during the OOD foraging.
EqVal UValues UItemNum USetSize
UnPre
GF 1.6 0.4 0.61 0.39
pre-GF 0.93 0.83 0.18 0.79
IVSN 76.03 47.6 47.05 77.76
IVSN-NN 72.72 46.85 47.18 64.03
Ours 81.63 70.87 65.16 72.34
Table S6: NormScore (%) of external baselines tested in OOD tasks. Best is in bold.
S3.9 Qualitative results of humans and our VF model
We visualized the scanpaths and click locations of our VF model and a human subject in Fig. S6B and C, respectively. As item positions were shuffled every three seconds in our experiment, we only depict clicks and fixations that occurred before the first shuffle. Both the human and the model primarily clicked on the highest-value targets (red balls), selecting them in 3 out of 6 clicks, indicating that target values strongly influenced their click decisions.
Figure S6: Qualitative results of our VF model evaluated under the UnValEqPre condition. (A) displays the targets along with their corresponding values for this example trial. (B) illustrates the model’s scanpaths and click locations, while (C) presents those of a human subject. Yellow dots indicate fixations, with connecting red lines representing visual scanpaths, and black numbers denoting fixation order. Red squares mark clicked items.
@@
First, we observed that humans occasionally clicked on items they were not directly fixating on, while VF assumes eye movements always align with the locations at which foraging decisions are made. Second, a strong priming effect was evident in humans, especially when target values were equal, showing the long-lasting influence of prior experiences on human decisions. Our VF currently lacks the ability to model such long-term dependencies, as it does not have a working memory integrating reinforcements from past actions into current decisions. Third, in hybrid foraging, humans actively compare fixated items with those in memory, a process known as memory search. Our VF assumes perfect memory search, where all targets are compared to the fixated item simultaneously. Fourth, fixation duration is another important aspect of human eye movement decisions. However, our VF model currently lacks the ability to capture fixation duration. Lastly, real-world environments may present additional challenges, such as target occlusions and physical constraints imposed by scene contexts. Extending the study of hybrid visual foraging beyond simplistic stimuli in controlled experimental settings remains an intriguing research direction.